# Speculative End-Turn Detector for Efficient Speech Chatbot Assistant

### v1

@@ 6.1 Binary classification
| Methods | Synthetic data | Real data | ||||||
| Pre. | Rec. | F1 | Acc. | Pre. | Rec. | F1 | Acc. | |
| VAP (Ekstedt & Skantze, 2022) | 48.4 | 48.5 | 48.3 | 57.2 | ||||
| GRU (Chung et al., 2014) | ||||||||
| Wav2vec 2.0 (Baevski et al., 2020) | ||||||||
In Table 4, we report the binary classification results of ETD models. From the table, we observe that the GRU models—which are much lighter than other models—achieve a significantly lower accuracy than other methods, with around 20%p gap in both synthetic and real datasets. This result motivates us to develop a computationally efficient framework that can achieve both efficiency and accuracy.
@@ 6.2 Real-time audio segmentation
| VAP (Ekstedt & Skantze, 2022) | 92.9 | 87.7 | 17.6 | 10.7 |
| GRU (Chung et al., 2014) | 85.5 | 76.2 | 24.8 | 14.7 |
| Wav2vec 2.0 (Baevski et al., 2020) | 94.7 | 90.2 | 30.3 | 17.9 |
| Ours | 94.0 | 88.9 | 28.0 | 16.4 |
In Table 5, we compare the real-time audio segmentation performances of the proposed SpeculativeETD against three baseline methods. The results demonstrate that the SpeculativeETD achieves a performance that is comparable to a larger-scale model, namely Wav2vec 2.0, less than 2%p difference in IoU. The proposed method also outperforms VAP, which requires much more computation.
@@ 6.3 Efficiency of SpeculativeETD
Methods Compute (MFLOPs) VAP 10,354.98 GRU 45.34 Wav2vec 2.0 34,971.68 SpeculativeETD 45.34 + 874.30 = 919.64 Table 6: FLOPs comparison for real-time processing.
@@
We evaluate the computational cost of our method in terms of FLOPs using 100 synthetic samples, as FLOPs for each sample are detailed in Table 6. By leveraging the server-side Wav2vec 2.0 model only when vital, SpeculativeETD achieves comparable performance to Wav2vec 2.0 with substantially lower FLOPs. Moreover, latency remains unaffected as Wav2vec 2.0 runs on a server GPU rather than on-device.
Methods Latency (ms) Load Init Execute Wav2vec 2.0 874.06 17.89 1500.32 GRU (SpeculativeETD) 1.16 3.85 0.26 Table 7: Latency analysis in a real-time on-device setting. The ‘Execute’ latency is the average inference time when the model performs inference every 100 ms on a 5-second speech audio.
@@
We verify whether the latency is suitable for real-time on-device deployment. We measure the latency via LiteRT 111https://ai.google.dev/edge/litert on iPhone mini 12 (iOS 16.3.1). We check model loading, initialization, and execution time, which is shown in Table 7. We averaged three runs of latency measurement, and the execution time was the average of 50 iterations of inference time when the model performed inference every 100 ms on a 5-second speech audio.
Our experiments demonstrate that SpeculativeETD, which utilizes GRU on the on-device side, achieves latency below 1 ms per 100 ms interval. As illustrated in Figure 5, the inference latency of Wav2vec 2.0 significantly increases with input length due to the computational overhead of transformer architectures, while GRU latency remains stable. These results confirm that SpeculativeETD is an excellent option for serving real-time on-device.
@@ 7 Conclusion
We introduce the ETD dataset, a novel publicly available dataset for end-turn detection in spoken dialogue systems, and propose SpeculativeETD, an efficient method for real-time end-turn detection. These contributions address critical challenges in developing more natural and responsive spoken dialogue agents, particularly those powered by large language models. The ETD dataset comprises synthetic and real data and is the first public dataset for end-turn detection research. This dataset enables more comprehensive training and evaluation of end-turn detection models by offering a diverse range of conversational scenarios, including artificially extended pauses and injected filler words. Also, real data further enhances the dataset’s utility in assessing model generalizability in real-world conversational settings. SpeculativeETD, represents a significant advancement in balancing efficiency and accuracy for real-time end-turn detection. Combining a lightweight GRU on-device model for on-device processing with a high-performance Wav2Vec 2.0 server-side model on the server improves accuracy while maintaining low latency. This two-stage framework effectively addresses the computational constraints of on-devices while leveraging the superior performance of larger models.
@@
All datasets used in this research are exclusively for academic purposes. We release open-source data according to original licensing terms and provide reproducible code for datasets that cannot be directly shared due to licensing constraints.
Appendix A More experiments
Full experiement result
| Methods | Speaking Unit | Gap | Pause | Average | ||||||||||
| Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | F1 Avg. | IoU Avg. | |
| VAP | ||||||||||||||
| GRU | ||||||||||||||
| Wav2vec 2.0 | ||||||||||||||
| Ours | ||||||||||||||
| Methods | Speaking Unit | Gap | Pause | Average | ||||||||||
| Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | Pre. | Rec. | F1 | IoU | F1 Avg. | IoU Avg. | |
| VAP | 51.6 | 35.1 | 41.8 | 26.4 | 2.0 | 6.5 | 3.1 | 1.6 | 10.3 | 6.2 | 7.8 | 4.0 | ||
| GRU | ||||||||||||||
| Wav2vec 2.0 | 55.7 | 23.5 | 33.03 | 19.78 | 21.1 | 54.1 | 30.34 | 17.88 | 28.7 | 26.4 | 27.48 | 15.93 | ||
| Ours | 50.3 | 19.4 | 28.04 | 16.30 | 20.3 | 53.4 | 29.44 | 17.26 | 27.4 | 25.7 | 26.56 | 15.31 | ||
Appendix B Data quality improvement
@@
We applied a structured segmentation process to better capture conversational dynamics. Consecutive speech fragments from the same speaker were merged into a Speaking Unit (SU) if separated by less than 200ms of silence. Silences longer than 200 ms were labeled Pause (if within the same speaker) or Gap (if between different speakers), based on established turn-taking criteria (Stivers et al., 2009; Heldner & Edlund, 2010). Using these labels, we grouped continuous SUs into segments, where each segment represents a stretch of speech bounded by either a Pause or a Gap. To focus on turn-relevant phenomena, we filtered out segments that did not contain any Pause or Gap, retaining only those that reflect actual hesitation or speaker transitions crucial for end-turn detection.
@@
To ensure the quality and consistency of the real-world speech data, we applied a language filtering step using the Whisper model (Radford et al., 2022). For each audio file, we used Whisper’s language detection capability to identify the spoken language. Since our focus is on English-based dialogue systems, we retained only the samples identified as English and discarded those in other languages. Table 10 summarizes the top 10 most frequent languages and the aggregate of other low-resource languages. Out of a total of 11,127 collected files, 8,022 were classified as English and included in the final dataset. We additionally conducted a human evaluation of language identification performance by Whisper, which is recognized well as English. We checked all the data, and the accuracy of identification was 99.07.
Appendix C Experiment details

### v2

@@ 6.1 Binary classification
In Table˜4, we report the binary classification results of ETD models. From the table, we observe that the GRU models—which are much lighter than other models—achieve a significantly lower accuracy than other methods, with around 20%p gap in both synthetic and real datasets. This result motivates us to develop a computationally efficient framework that can achieve both efficiency and accuracy.
@@ 6.2 Real-time audio segmentation
| VAP (Ekstedt and Skantze, 2022) | 90.6 | 84.8 | 33.2 | 25.9 |
| GRU (Chung et al., 2014) | 58.0 | 52.2 | 34.2 | 31.7 |
| Wav2Vec2 (Baevski et al., 2020) | 94.7 | 90.2 | 58.4 | 46.2 |
| SpeculativeETD (Ours) | 94.0 | 88.9 | 45.6 | 37.8 |
In Table˜5, we compare the real-time audio segmentation performances of the proposed SpeculativeETD against three baseline methods. The results demonstrate that the SpeculativeETD achieves a performance that is comparable to a larger-scale model, namely Wav2vec 2.0, on the synthetic set. The proposed method also outperforms VAP, which requires much more computation.
| Methods | Compute (MFLOPs) |
| VAP | 10,354.98 |
| GRU | 45.34 |
| wav2vec 2.0 | 34,971.68 |
| SpeculativeETD | 45.34 + 874.30 = 919.64 |
@@ 6.3 Efficiency of SpeculativeETD
We evaluate the computational cost of our method in terms of FLOPs using 100 synthetic samples. By leveraging the server-side Wav2vec 2.0 model only when vital, SpeculativeETD achieves comparable performance to Wav2vec 2.0 with substantially lower FLOPs. Moreover, latency remains unaffected as Wav2vec 2.0 runs on a server GPU rather than on-device.
@@
We verify whether the latency is suitable for real-time on-device deployment. We measure the latency via LiteRT111https://ai.google.dev/edge/litert on iPhone mini 12 (iOS 16.3.1). We check model loading, initialization, and execution time, which is shown in Table˜7. We averaged three runs of latency measurement, and the execution time was the average of 50 iterations of inference time when the model performed inference every 100 ms on a 5-second speech audio.
Our experiments demonstrate that SpeculativeETD, which utilizes GRU on the on-device side, achieves latency below 1 ms per 100 ms interval. As illustrated in Fig.˜5, the inference latency of Wav2vec 2.0 significantly increases with input length due to the computational overhead of transformer architectures, while GRU latency remains stable. On top of on-device inference, we also measured end-to-end audio-transfer latency over 5G and Wi-Fi; round-trip times stay within – across payload sizes up to , well inside the turn-taking threshold (see Appendix˜E). These results confirm that SpeculativeETD is an excellent option for serving real-time on-device.
| Methods | Latency (ms) | ||
| Load | Init | Execute | |
| wav2vec 2.0 | 874.06 | 17.89 | 1500.32 |
| GRU (SpeculativeETD) | 1.16 | 3.85 | 0.26 |
@@ 7 Conclusion
We introduce the OpenETD dataset, a novel publicly available dataset for end-turn detection in spoken dialogue systems, and propose SpeculativeETD, an efficient method for real-time end-turn detection. These contributions address critical challenges in developing more natural and responsive spoken dialogue agents, particularly those powered by large language models. The OpenETD dataset comprises synthetic and real data and is the first public dataset for end-turn detection research. This dataset enables more comprehensive training and evaluation of end-turn detection models by offering a diverse range of conversational scenarios, including artificially extended pauses and injected filler words. Also, real data further enhances the dataset’s utility in assessing model generalizability in real-world conversational settings. SpeculativeETD, represents a significant advancement in balancing efficiency and accuracy for real-time end-turn detection. Combining a lightweight GRU on-device model for on-device processing with a high-performance Wav2Vec 2.0 server-side model on the server improves accuracy while maintaining low latency. This two-stage framework effectively addresses the computational constraints of on-devices while leveraging the superior performance of larger models.
Limitations
Our work has several limitations. First, the OpenETD dataset primarily focuses on English conversations, which may limit its applicability to other languages with different turn-taking patterns. Second, while SpeculativeETD demonstrates strong performance, the reliance on server-side computation for Gap/Pause classification introduces network latency considerations not addressed in our current evaluation. Third, our synthetic data generation, while effective, may not fully capture the complexity of spontaneous human speech patterns.
@@
All datasets used in this research are used exclusively for academic purposes. We release open-source processing code and, for data that cannot be directly shared due to licensing constraints, provide reproducible download scripts. We build on four data sources with distinct terms: MultiWOZ (Budzianowski et al., 2018) (Apache 2.0), Google Cloud Text-to-Speech (used under the Google Cloud Terms of Service), YouTube audio (redistributed only as URLs with download scripts; we do not rehost audio files), and the Buckeye Corpus (Pitt et al., 2005) (used under its Academic License). Full licensing details are provided in Appendix˜F.
Acknowledgments
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (Nos. RS-2024-00457882, RS-2019-II191906, 2022-0-00713) and in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00453301, RS-2025-24873016, RS-2026-25494004).
Appendix A Synthetic pause generation: statistical validation
Our synthetic pipeline does not insert pauses at random positions. Instead, we detect micro-hesitations naturally produced by the TTS model (Braunschweiler and Buchholz, 2011; Hwang et al., 2023; Yang et al., 2024) and extend their duration using an Erlang distribution fitted to real pause durations.
Duration.
Table˜8 reports Kolmogorov–Smirnov (KS) two-sample statistics and Cohen’s comparing synthetic and real gap/pause durations. The KS statistic for gap duration is only (Cohen’s ), indicating that the Erlang-fitted generator closely matches real duration distributions.
Position.
In real conversational data, both pause and gap positions within an utterance are close to uniform (KS vs. Uniform: pause , gap ), reflecting the inherent unpredictability of pause placement. Synthetic positions show a larger divergence (KS for both pause and gap). We acknowledge this as a limitation; however, exact position replication is inherently difficult given that real positions themselves are near-uniform.
Augmentation value.
The synthetic data therefore functions as augmentation rather than exact replication. Table˜9 shows that training SpeculativeETD on the mixture consistently outperforms training on either split alone, confirming the practical value of the synthetic set.
| Comparison | KS stat | Cohen’s |
| Duration | ||
| Syn vs. Real gap duration | 0.083 | 0.12 |
| Syn vs. Real pause duration | 0.092 | 0.15 |
| Position within utterance | ||
| Real pause vs. Uniform | 0.091 | – |
| Real gap vs. Uniform | 0.094 | – |
| Syn vs. Real gap (position) | 0.238 | 0.23 |
| Syn vs. Real pause (position) | 0.238 | 0.27 |
| Training mode | Real F1 | Real IoU | F1 | IoU |
| Mix (syn + real) | 45.6 | 37.8 | – | – |
| Real only | 43.1 | 36.3 | ||
| Synthetic only | 44.0 | 36.7 |
Appendix B On-device GRU input specification
We detail the input pipeline of the on-device GRU used in SpeculativeETD.
Mel-spectrogram input.
Each input chunk corresponds to of audio sampled at . We extract a -dim log–mel spectrogram with a window and a hop, yielding frames per chunk, followed by per-utterance z-score normalization. The resulting chunk tensor has shape .
Conv2D frontend.
Two Conv2D layers downsample the chunk: , producing a -dim per-chunk feature vector after flattening.
GRU backbone.
A single-layer GRU (hidden size ) consumes the chunk features autoregressively, and a linear head outputs the binary (SU vs. non-SU) logits. The entire on-device model has approximately parameters.
Filler word vocabulary.
During synthetic data generation (Section˜3.1), we inject filler words drawn from the following vocabulary: “uh”, “um”, “ah”, “er”, “hmm”, “mhm”, “uh huh”, “like”, “you know”, “so”, “actually um”, and “basically um”. These cover the common English hesitation markers observed in conversational speech.
Appendix C Human validation of pause/gap labels
To quantify label noise introduced by the automatic diarization + 200 ms rule, we performed a human validation study. Three annotators each judged randomly sampled real clips ( Pause, Gap). Each clip was presented with of audio before the silence and after, and annotators reported (i) whether the automatic label was correct and (ii) the diarization segmentation quality on a – scale.
Table˜10 summarizes the results. Pause labels are highly reliable (human–auto agreement ), while Gap labels are moderately noisier (76.1%) because speaker-change boundaries are intrinsically harder to localize. Overall diarization quality averages , supporting the use of the automatic pipeline. The threshold follows the widely adopted convention in turn-taking research (Stivers et al., 2009; Heldner and Edlund, 2010).
| Metric | Pause () | Gap () | Overall () |
| Human–Auto agreement | 94.0% | 76.1% | 85.4% |
| Diarization quality (1–5) |
Appendix D Domain analysis: synthetic vs. real
To characterize the mismatch between synthetic and real conversations, we analyzed audio-level and speaker-level statistics (Table˜11). Signal-to-noise ratio (SNR) is computed from the raw waveform, Word-Per-Second (WPS) from Whisper-tiny transcripts, gender is derived from metadata, and speaker count reflects the distinct voice identities present.
The two largest domain gaps are SNR ( vs. ) and speaking-rate variability ( vs. ). Real data also exhibits a richer emotion distribution (with a substantially higher share of sad speech) and spans five English accents, whereas synthetic data uses only two TTS voices (EN-US and EN-BR); see Table˜11.
| Dimension | Synthetic | Real |
| Signal-to-noise ratio (dB) | ||
| Word per second | ||
| Gender (Female %) | 66.2 | 39.4 |
| Speaker count | 3 TTS voices | 200 |
| Emotion (%) | Accent, Real (%) | |||||||||
| Happy | Sad | Angry | Disgust | Other | US | England | Canada | Indian | Australia | |
| Real | 33 | 33 | 11 | 5 | 18 | 44.0 | 16.6 | 16.6 | 14.6 | 8.2 |
| Synthetic | 30 | 8 | 19 | 30 | 13 | Two TTS voices (EN-US, EN-BR) | ||||
Appendix E End-to-end audio transfer latency
Although SpeculativeETD operates asynchronously—the on-device GRU continues detecting silence while the server classifies in the background—excessively long transfer times could still delay the classification decision. We therefore measured real audio-transfer round-trip time (RTT) from an iPhone 12 mini to a remote server via WebSocket, without model inference, across six payload sizes (– of audio; trials each).
Table˜12 reports the results. On 5G, RTT ranges from to ; on Wi-Fi, to —both comfortably inside the turn-taking threshold. RTT is dominated by fixed network latency rather than payload size; a payload increase () adds only about on 5G.
| Audio | Payload | Mean RTT | P50 | P95 |
| 5G network | ||||
| 100 ms | 3.1 KB | 106.0 ms | 104.0 ms | 115.0 ms |
| 500 ms | 15.6 KB | 106.4 ms | 104.0 ms | 119.0 ms |
| 1 s | 31.3 KB | 106.4 ms | 104.0 ms | 119.0 ms |
| 3 s | 93.8 KB | 107.2 ms | 103.0 ms | 110.0 ms |
| 6 s | 187.5 KB | 109.9 ms | 108.0 ms | 128.0 ms |
| 10 s | 312.5 KB | 116.2 ms | 114.0 ms | 129.0 ms |
| Wi-Fi | ||||
| 100 ms | 3.1 KB | 100.2 ms | 95.0 ms | 131.0 ms |
| 500 ms | 15.6 KB | 101.6 ms | 94.0 ms | 142.0 ms |
| 1 s | 31.3 KB | 107.9 ms | 99.0 ms | 178.0 ms |
| 3 s | 93.8 KB | 98.8 ms | 94.0 ms | 121.0 ms |
| 6 s | 187.5 KB | 107.0 ms | 100.0 ms | 128.0 ms |
| 10 s | 312.5 KB | 139.7 ms | 115.0 ms | 225.0 ms |
Appendix F Licensing of data sources
Table˜13 summarizes the licenses of the data sources used to build OpenETD. We release the dataset and processing code in a manner consistent with each underlying license: the MultiWOZ-derived synthetic data inherits Apache 2.0; YouTube audio is not redistributed (we share URLs and download scripts only); Buckeye Corpus is used under its Academic License.
| Dataset / Source | License |
| MultiWOZ (Budzianowski et al., 2018) | Apache 2.0 |
| Google Cloud Text-to-Speech | Google Cloud Terms of Service |
| YouTube audio | URLs + download scripts (not redistributed) |
| Buckeye Corpus (Pitt et al., 2005) | Academic License |
Appendix G Data quality improvement
@@
We applied a structured segmentation process to better capture conversational dynamics. Consecutive speech fragments from the same speaker were merged into a Speaking Unit (SU) if separated by less than 200 ms of silence. Silences longer than were labeled Pause (if within the same speaker) or Gap (if between different speakers), based on established turn-taking criteria (Stivers et al., 2009; Heldner and Edlund, 2010). Using these labels, we grouped continuous SUs into segments, where each segment represents a stretch of speech bounded by either a Pause or a Gap. To focus on turn-relevant phenomena, we filtered out segments that did not contain any Pause or Gap, retaining only those that reflect actual hesitation or speaker transitions crucial for end-turn detection.
@@
To ensure the quality and consistency of the real-world speech data, we applied a language filtering step using the Whisper model (Radford et al., 2022). For each audio file, we used Whisper’s language detection capability to identify the spoken language. Since our focus is on English-based dialogue systems, we retained only the samples identified as English and discarded those in other languages. Table˜14 summarizes the top 10 most frequent languages and the aggregate of other low-resource languages. Out of a total of 11,127 collected files, 8,022 were classified as English and included in the final dataset. We additionally conducted a human evaluation of language identification performance by Whisper, which is recognized well as English. We checked all the data, and the accuracy of identification was .
Appendix H Experiment details